# Optimal Monotone Depth-Three Circuit Lower Bounds for Majority

### v1

@@ 3.1 Cliques
For $l\geq k\geq 1$, let ${\mathbb{K}}_{l}^{k}$ denote the $k$-system of all $k$-sets on $l$ elements.
The following two cliques are widely used in our constructions: ${\mathbb{K}}_{3}^{3}=\{\{1,2,3\}\}$ is a $1$-threshold 33-system where $\theta_{1}({\mathbb{K}}_{3}^{3})=3$; ${\mathbb{K}}_{4}^{3}=\{\{1,2,3\},\{1,2,4\},\{1,3,4\},\{2,3,4\}\}$ is a $2$-threshold 33-system where $\theta_{2}({\mathbb{K}}_{4}^{3})=6$.
@@
${\mathbb{K}}_{l}^{k}$ is an extremal (lk+1)(l-k+1)-threshold $k$-uniform set system on $l$ elements where $n({\mathbb{K}}_{l}^{k})=l,$ $m({\mathbb{K}}_{l}^{k})=\binom{l}{k}$, $\tau({\mathbb{K}}_{l}^{k})=l-k+1,$ and $\Theta(l,l-k+1,k)=\theta_{l-k+1}({\mathbb{K}}_{l}^{k})=\binom{l}{l-k+1}$.
@@ 3.2 Turán-style constructions
Let $n\geq 3$. Constructions from here on are primarily limited to $k=3$. Consider the following (n,n3,3)(n,n-3,3)-Turán system ${\mathbb{T}}_{n}^{3}$ which is obtained by what we call a TuránTur\'{a}n-style construction [16]. Partition $n$ into three parts where the parts have sizes as equal as possible and organize them in a circular fashion. ${\mathbb{T}}_{n}^{3}$ consists of the following $3$-sets:
@@
    For each part, all 3-sets of the part
@@
    For each part, all 3-sets of the form where two elements come from the part and one element from the adjacent part in the circular order.
Conjecture 3.3 (Turán).
[…]
The conjecture implies
Fact 3.4.
Conjecture 3.3 implies
[…]
@@
${\mathbb{T}}_{5}^{3}=\{\{1,2,3\},\{3,4,5\},\{1,2,4\}\}$ is a $2$-threshold 33-uniform set system, $n({\mathbb{T}}_{5}^{3})=5$, $m({\mathbb{T}}_{5}^{3})=3$, $\tau({\mathbb{T}}_{5}^{3})=2$, and $\theta_{2}({\mathbb{T}}_{5}^{3})=7$.
@@
Proofs of the following statement are provided in the appendix (Theorems A.1 and A.2).
Fact 3.7.
@@ 3.3 Constructions for $0\leq t\leq\lfloor n/k\rfloor$ and $n-k+1\leq t\leq n$
In this section, we provide constructions that work for all $n,k\geq 1$, $0\leq t\leq\lfloor n/k\rfloor$ and $n-k+1\leq t\leq n$. We need the following notation to represent larger set systems composed of smaller set systems. If $\mathbb{F}_{1}=(V_{1},F_{1})$ and $\mathbb{F}_{2}=(V_{2},F_{2})$ are $k$-systems over disjoint sets of elements, the sum $\mathbb{F}_{1}+\mathbb{F}_{2}$ of $\mathbb{F}_{1}$ and $\mathbb{F}_{2}$ represents the $k$-system $\mathbb{F}=(V_{1}\cup V_{2},F_{1}\cup F_{2})$. For a $k$-system $\mathbb{F}$, $\ell\cdot\mathbb{F}$ represents the $k$-system obtained by summing $\ell$ disjoint copies of $\mathbb{F}$.
Let $\mathbb{S}^{k}_{n,t}=t\cdot{\mathbb{K}}_{k}^{k}$. For $n\geq 1$ and $0\leq t\leq\lfloor n/k\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique optimal construction for any $k\geq 1$.
@@
For $k,n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{k}\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique (up to isomorphism) extremal $t$-threshold $k$-system on $n$ elements.
@@
The (ki+1)(k-i+1)-uniform set system consisting of all (ki+1)(k-i+1)-sets is an extremal $t$-threshold $k$-system and $\Theta(n,n-k+i,k)=\binom{n}{n-k+i}$.
Our constructions for $n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{2}\rceil$ are limited to the case $k=3$. They use ${\mathbb{K}}_{3}^{3}$, ${\mathbb{K}}_{4}^{3}$, ${\mathbb{T}}_{5}^{3}$, and ${\mathbb{T}}_{6}^{3}$ as building blocks. The following important reparameterization is useful to express our constructions.
Definition 3.10.
For $k=3$, we rewrite n=3tsn=3t-s where $s$ is called the deficit of the elements (variables or vertices) with respect to the threshold $t$. To emphasize the role of $s$, we specialize $\Theta(n,t,k)$ for $k=3$ and use the function $\Phi_{0}(s,t)$ to denote the maximum number of $t$-transversals of a $t$-threshold 33-system with deficit $s$. In other words, $\Phi_{0}(s,t)=\Theta(3t-s,t,3)$.
Let $s\leq t$ and $t\geq 0$ (equivalently $n\geq 2t\geq 0$). Let $\mathbb{P}^{3}_{s,t}$ be the $3$-system on n=3tsn=3t-s elements defined as follows.
[…]
Fact 3.11.
$\mathbb{P}^{3}_{s,t}$ is a $t$-threshold 33-uniform set system.
Fact 3.12.
@@ 3.4 Constructions for $n\geq 1,k=3$ and $0\leq t\leq\lfloor\frac{n}{2}\rfloor$
${\mathbb{K}}_{3}^{3}+{\mathbb{T}}_{6}^{3}$ can be replaced ${\mathbb{K}}_{4}^{3}+{\mathbb{T}}_{5}^{3}$ with out affecting the number of minimum-size transversals.
We summarize the results of this section by the following theorem which provides lower bounds for the maximum number of minimum-size transversals $\Phi(s,t)$ of $3$-systems over n=3tsn=3t-s elements with transversal number and deficit satisfying the conditions, $s\leq t$ and $t\geq 0$.
@@
The first equation follows from [4]. The second equation is straightforward. The remaining two inequalities follow from Fact 3.12. ∎
@@ 4 Upper Bounds on $\Phi_{0}(s,t)$
In this section, we establish matching upper bounds on the number of transversals $\Phi_{0}(s,t)$ of extremal $t$-threshold 33-systems for $t\geq 0$ and $s\leq t$. We prove these bounds by induction on 3ts3t-s and $t$. As we do induction on $s$ and $t$, we have to deal with restricted set systems which motivates the following definitions.
For $i\in\{1,2,3,4\}$, we say a $3$-system $\mathbb{F}$ is a type $i$ 33-system if it contains at least $i$ 22-sets. We further specialize the type 2 systems: $\mathbb{F}$ is a type $2o$ 33-system if it contains two $2$-sets which have exactly one element in common and a type $2d$ 33-system if it contains two $2$-sets which are disjoint. We also restrict type 3 systems as follows: $\mathbb{F}$ is a type $3x$ 33-system if it contains three $2$-sets which do not have an element in common. By extension, unrestricted $3$-systems are of type 0. For $i\in\{1,2d,2o,3x,4\}$, let
  • ${\Xi}_{i}$ denote the family of $3$-system of type $i$,
  • ${\Xi}_{i}(s,t)$ denote the family of $t$-threshold 33-system of type $i$ over n=3tsn=3t-s elements, and
  • $\Phi_{i}(s,t)$ denote the maximum number of $t$-transversals over all type $i$ $t$-threshold 33-systems over 3ts3t-s elements.
@@
3-systems with at least one $2$-set
@@
3-systems with two $2$-sets which overlap in one element
@@
3-systems with two $2$-sets which are disjoint
@@
3-systems with three $2$-sets which do not have an element in common
@@
3-systems with four $2$-sets
@@ 4.1 Outline of the proof
The proof assumes that $t\geq 1$ as the theorem is trivially true otherwise. We divide the task of proving the upper bounds into six parts. In part $i\in\{0,1,2d,2o,3x,4\}$, we prove the upper bounds for $\Phi_{i}(s,t)$, the number of $t$-transversals for set systems in ${\Xi}_{i}(s,t)$ assuming the upper bounds for $\Phi_{j}(s,t)$ for smaller values of 3ts3t-s or $t$ for $j\in\{0,1,2d,2o,3x,4\}$.
@@
Our induction step has the following pattern: For each $i\in\{0,1,2d,2o,3x,4\}$, we develop a set of properties $\mathscr{P}(i,j)$ such that each $\mathbb{F}\in{\Xi}_{i}(s,t)$ satisfies one of the properties. Each property $\mathscr{P}(i,j)$ identifies a minimal collection of sets included in $\mathbb{F}$ which we call an anchor.
@@
For example, we have a set of six properties $\mathscr{P}(0,j)$ ($1\leq j\leq 5$) such that every set system $\mathbb{F}$ in ${\Xi}_{0}$ satisfies one of the six properties. $\mathscr{P}(0,1)$ states that there exists a pair of elements that appears in three distinct sets. If $\mathbb{F}\in{\Xi}_{0}$ satisfies this property, we know that $\mathbb{F}$ includes the anchor $\mathbb{A}=\{\{a,b,c\},\{a,b,d\},\{a,b,e\}\}$ for some distinct elements $a,b,c,d,$ and $e$.
@@
Since every $t$-transversal of $\mathbb{F}$ is an extension of some transversal of the anchor, we enumerate the transversals of the anchor and for each such transversal we bound the number of extensions that lead to $t$-transversals. For enumerating the transversals of the anchor, We select certain elements (which we call core elements) that appear in the anchor and enumerate all transversals of the anchor based on the inclusion/exclusion of the core elements and any elements which are forced as a consequence. Let $\gamma$ be one such transversal and assume that it includes $j$ elements and excludes $e$ elements. The transversal $\gamma$ induces a restricted system $\mathbb{F}^{\prime}$ of sets. Specifically, $\mathbb{F}^{\prime}$ is obtained by eliminating all sets from $\mathbb{F}$ containing any included element of $\gamma$ and by deleting any excluded element from each set of $\mathbb{F}$. It is clear that $\mathbb{F}^{\prime}$ is a (tj)(t-j)-threshold 33-system over njen-j-e elements and its deficit is 3tn2j+e3t-n-2j+e. We also ensure that $\mathbb{F}^{\prime}\in{\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3x,4\}$. We now use induction to bound the number of (tj)(t-j)-transversals of $\mathbb{F}^{\prime}$.
There is another wrinkle to the induction step. While our discussion focused on a minimal anchor, the actual choice of the anchor depends on the occurrence pattern of core elements in sets outside of the anchor. We thus deal with several cases and in each case we work with an anchor which is obtained by extending a minimal anchor.
To continue with our example of the property $\mathscr{P}(0,1)$, we select $a$ and $b$ as our core variables. If there is an $x\in\{a,b\}$ such that whenever it appears in a set of $\mathbb{F}$ the other element, $\{a,b\}-\{x\}$ also appears in the set, we stick with the anchor $\mathbb{A}$. If not, we extend the anchor $\mathbb{A}$ to $\mathbb{A}^{\prime}$ include the additional sets in which $a$ or $b$ appear by themselves without the other element.
In each case, we enumerate all transversals of the anchor by including/excluding the core variables. The transversals (specified by the inclusion/exclusion of core elements), the resulting restricted set systems of $\mathbb{F}$ together with their parameters, and the inductive bounds of the number of transversals of the restricted systems are represented in a table as shown below for the anchor $\mathbb{A}$ when there is an $x\in\{a,b\}$ such that whenever it appears in a set of $\mathbb{F}$ the other element, $\{a,b\}-\{x\}$ also appears in the set. Without loss of generality (due to symmetry between $a$ and $b$), we assumed that $a$ does not appear in any set of $\mathbb{F}$ in which $b$ does not appear.
@@
The columns with the label Core elements contain entries indicating whether the core elements are included in or excluded from the transversals of the anchor, $1$ signals inclusion and 0 signals exclusion. The column labeled Forced elements lists elements which are forced to be included or excluded as a consequence of the inclusion/exclusion of the core elements. The columns under the label Subsystem parameters describe the values of the parameters $n^{\prime}$, $s^{\prime}$, and $t^{\prime}$ as well as the type of the resulting restricted system. Each row (except the last row) represents a transversal of the anchorḞor each such row, the columns labeled Inductive bound contains the inductive upper bounds on the number of $t^{\prime}$-transversals for the restricted subsystem expressed as a fraction of the claimed upper bound on the number of $t$-transversals for $\mathbb{F}$ depending on the parity of $s$. In our example, we bound the number of transversals as a fraction of $(\frac{2}{3})^{\frac{s}{2}}3^{t}$ for even $s$. For odd $s$, we bound the number of transversals as a fraction of $(\frac{7}{9})(\frac{2}{3})^{\frac{(s-1)}{2}}3^{t}$. The last row contains the sum of the fractions for even and odd parities of $s$. The induction step is successful if the fraction add up to less than or equal to $1$.
Core elements Forced elements Subsystem parameters Inductive bound

### v2

@@ 3.1 Cliques
For $l\geq k\geq 1$, let ${\mathbb{K}}_{l}^{k}$ denote the monotone $k$-CNF of all monotone $k$-clauses on $l$ variables.
The following two cliques are widely used in our constructions: ${\mathbb{K}}_{3}^{3}=\{\{1,2,3\}\}$ is a $1$-threshold monotone $3$-CNF where $\theta_{1}({\mathbb{K}}_{3}^{3})=3$; ${\mathbb{K}}_{4}^{3}=\{\{1,2,3\},\{1,2,4\},\{1,3,4\},\{2,3,4\}\}$ is a $2$-threshold monotone $3$-CNF where $\theta_{2}({\mathbb{K}}_{4}^{3})=6$.
@@
${\mathbb{K}}_{l}^{k}$ is an extremal (lk+1)(l-k+1)-threshold $k$-uniform monotone CNF on $l$ variables where $n({\mathbb{K}}_{l}^{k})=l,$ $m({\mathbb{K}}_{l}^{k})=\binom{l}{k}$, $\tau({\mathbb{K}}_{l}^{k})=l-k+1,$ and $\Theta(l,l-k+1,k)=\theta_{l-k+1}({\mathbb{K}}_{l}^{k})=\binom{l}{l-k+1}$.
@@ 3.2 Turán-style constructions
Let $n\geq 3$. We observe that $\Theta(n,t,3)$ is closely related to the hypergraph Turán problem. In particular, t=n3t=n-3 exactly captures the tetrahedron problem [15, 9]. Consider the following (n3)(n-3)-threshold formula ${\mathbb{T}}_{n}^{3}$ which is obtained by what we call a TuránTur\'{a}n-style construction. Partition the $n$ variables into three parts where the parts have sizes as equal as possible and organize them in a circular fashion. ${\mathbb{T}}_{n}^{3}$ consists of the following $3$-clauses:
@@
    For each part, all 3-clauses of the part
@@
    For each part, all 3-clauses of the form where two variables come from the part and one variable from the adjacent part in the circular order.
@@
${\mathbb{T}}_{5}^{3}=\{\{1,2,3\},\{3,4,5\},\{1,2,4\}\}$ is a $2$-threshold 33-uniform monotone CNF, $n({\mathbb{T}}_{5}^{3})=5$, $m({\mathbb{T}}_{5}^{3})=3$, $\tau({\mathbb{T}}_{5}^{3})=2$, and $\theta_{2}({\mathbb{T}}_{5}^{3})=7$.
@@
Proofs of the following statement are provided in the appendix (Theorems A.2 and A.3).
Fact 3.5.
@@ 3.3 Constructions for $0\leq t\leq\lfloor n/k\rfloor$ and $n-k+1\leq t\leq n$
In this section, we provide constructions that work for all $n,k\geq 1$, $0\leq t\leq\lfloor n/k\rfloor$ and $n-k+1\leq t\leq n$. We need the following notation to represent larger CNFs composed of smaller CNFs. If $\mathbb{F}_{1}=(V_{1},F_{1})$ and $\mathbb{F}_{2}=(V_{2},F_{2})$ are monotone $k$-CNFs over disjoint sets of variables, the sum $\mathbb{F}_{1}+\mathbb{F}_{2}$ of $\mathbb{F}_{1}$ and $\mathbb{F}_{2}$ represents the monotone $k$-CNF $\mathbb{F}=(V_{1}\cup V_{2},F_{1}\cup F_{2})$. For a monotone $k$-CNF $\mathbb{F}$, $\ell\cdot\mathbb{F}$ represents the monotone $k$-CNF obtained by summing $\ell$ disjoint copies of $\mathbb{F}$.
Let $\mathbb{S}^{k}_{n,t}=t\cdot{\mathbb{K}}_{k}^{k}$. For $n\geq 1$ and $0\leq t\leq\lfloor n/k\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique extremal construction for any $k\geq 1$.
@@
For $k,n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{k}\rfloor$, $\mathbb{S}^{k}_{n,t}$ is the unique (up to isomorphism) extremal $t$-threshold monotone $k$-CNF on $n$ variables.
@@
Fact 3.7.
The (ki+1)(k-i+1)-uniform monotone CNF consisting of all (ki+1)(k-i+1)-clauses is an extremal $t$-threshold monotone $k$-CNF and $\Theta(n,n-k+i,k)=\binom{n}{n-k+i}$.
Our constructions for $n\geq 1$ and $0\leq t\leq\lfloor\frac{n}{2}\rceil$ are limited to the case $k=3$. They use ${\mathbb{K}}_{3}^{3}$, ${\mathbb{K}}_{4}^{3}$, ${\mathbb{T}}_{5}^{3}$, and ${\mathbb{T}}_{6}^{3}$ as building blocks. Our constructions vary based on the parity of 3tn3t-n, so the following reparameterization is useful.
Definition 3.8.
For $k=3$, we rewrite n=3tsn=3t-s where $s$ is called the deficit of the variables (elements or vertices) with respect to the threshold $t$. Our constructions and bounds depend on the parity of $s$. To emphasize the role of $s$, we specialize $\Theta(n,t,k)$ for $k=3$ and use the function $\Phi_{0}(s,t)$ to denote the maximum number of $t$-transversals of a $t$-threshold monotone $3$-CNF with deficit $s$. In other words, $\Phi_{0}(s,t)=\Theta(3t-s,t,3)$.
Let $s\leq t$ and $t\geq 0$ (equivalently $n\geq 2t\geq 0$). Let $\mathbb{P}^{3}_{s,t}$ be the $3$-CNF over n=3tsn=3t-s variables defined as follows.
[…]
@@ 3.4 Constructions for $n\geq 1,k=3$ and $0\leq t\leq\lfloor\frac{n}{2}\rfloor$
$\mathbb{P}^{3}_{s,t}$ is a $t$-threshold 33-uniform monotone CNF.
Fact 3.10.
@@
${\mathbb{K}}_{3}^{3}+{\mathbb{T}}_{6}^{3}$ can be replaced with ${\mathbb{K}}_{4}^{3}+{\mathbb{T}}_{5}^{3}$ without affecting the number of minimum-size transversals.
We summarize the results of this section by the following theorem. It provides lower bounds for the maximum number of minimum-size transversals $\Phi_{0}(s,t)$ of monotone $3$-CNFs over n=3tsn=3t-s elements with transversal number and deficit satisfying the conditions $s\leq t$ and $t\geq 0$.
@@
The first equation follows from [4]. The second equation is straightforward. The remaining two inequalities follow from Fact 3.10. ∎
@@ 4 Upper Bounds on $\Phi_{0}(s,t)$
In this section, we outline our plan to establish matching upper bounds on the number of transversals $\Phi_{0}(s,t)$ of extremal $t$-threshold monotone $3$-CNFs for $t\geq 0$ and $s\leq t$. We prove these bounds by induction on 3ts3t-s and $t$. As we do induction on $n$ and $t$, we have to deal with restricted monotone CNFs, which motivates the following definitions.
For $i\in\{1,2,3,4\}$, we say a monotone $3$-CNF $\mathbb{F}$ is a type $i$ monotone $3$-CNF if it contains at least $i$ monotone $2$-clauses. We further specialize the type 2 systems: $\mathbb{F}$ is a type $2o$ monotone $3$-CNF if it contains two monotone $2$-clauses which have exactly one variable in common and a type $2d$ 33-CNF if it contains two monotone $2$-clauses which are disjoint. As for type 3 CNFs, we only need to work with a restricted class of type 3 formulas, that is, formulas that do not contain three monotone $2$-clauses which have a variable in common. Since we do not consider other subtypes of type 3 CNFs, we will use type 3 CNFs to refer to the restricted class. By extension, unrestricted monotone $3$-CNFs are of type 0. For $i\in\{1,2d,2o,3,4\}$, let
  • ${\Xi}_{i}$ denote the family of monotone $3$-CNFs of type $i$,
  • ${\Xi}_{i}(s,t)$ denote the family of $t$-threshold monotone $3$-CNFs of type $i$ over n=3tsn=3t-s variables, and
  • For $i\in\{0,1\}$, let $\Phi_{i}(s,t)$ denote the maximum number of $t$-transversals over all type $i$ $t$-threshold monotone $3$-CNFs over 3ts3t-s variables. For $i\in\{2d,2o,3,4\}$, let $\Phi_{i}(s,t)$ denote the maximum number of $t$-transversals over all type $i$ $t$-threshold monotone $3$-CNFs $F$ over 3ts3t-s variables where there is no pair of variables that appear in three clauses of $F$.
@@
Monotone $3$-CNFs with at least one $2$-clause
@@
Monotone $3$-CNFs with two $2$-clauses which overlap in one variable
@@
Monotone $3$-CNFs with two $2$-clauses which are disjoint
@@
Monotone $3$-CNFs with three $2$-clauses which do not have an variable in common
@@
Monotone $3$-CNFs with four $2$-clauses
@@
Recall that we are primarily interested in the number of transversals of size $\lfloor n/2\rfloor$ in an $n$-vertex 3-graph with no transversal of smaller size. By setting $t=\lfloor n/2\rfloor$ and $s=t$, our bound follows.
Corollary 4.2.
$\Theta(n,\lfloor n/2\rfloor,3)\leq 6^{n/4}$.
@@ 4.1 Outline of the proof
The theorem trivially holds for $t=0$, so our proof focuses on the inductive step. For each $i$ and for all $j\in\{0,1,2d,2o,3,4\}$, we assume our upper bounds on $\Phi_{j}(s,t)$, the number of $t$-transversals of monotone CNFs in ${\Xi}_{j}(s,t)$, hold for all smaller values of 3ts3t-s or $t$. We use this to prove the upper bounds also hold for $\Phi_{i}(s,t)$.
@@
Our inductive step has the following pattern: For each $i\in\{0,1,2d,2o,3,4\}$, we develop a set of properties $\mathscr{P}(i,j)$ such that each $\mathbb{F}\in{\Xi}_{i}(s,t)$ satisfies at least one of the properties. Each property $\mathscr{P}(i,j)$ identifies a minimal collection of clauses included in $\mathbb{F}$, which we call an anchor.
@@
For example, we have a set of six properties $\mathscr{P}(0,j)$ ($1\leq j\leq 6$) such that every monotone CNF $\mathbb{F}$ in ${\Xi}_{0}$ satisfies at least one of the six properties. $\mathscr{P}(0,1)$ states that there exists a pair of variables that appears in three distinct clauses. If $\mathbb{F}\in{\Xi}_{0}$ satisfies this property, we know that $\mathbb{F}$ includes the clauses in $\mathbb{A}=\{\{a,b,c\},\{a,b,d\},\{a,b,e\}\}$ where the variables $a,b,c,d,$ and $e$ are distinct. We consider two cases to determine the anchor: either $a$ always appears together with $b$ in $\mathbb{F}$ or not. In the first case, we take $\mathbb{A}$ as our anchor. In the other case, our anchor will be $\mathbb{A}^{\prime}=\mathbb{A}\cup\{C\}$ where $C\in\mathbb{F}$ is a clause that contains $a$ but not $b$.
@@
Since every $t$-transversal of $\mathbb{F}$ is an extension of some transversal of the anchor, we enumerate all transversals of the anchor, and for each such transversal we bound the number of extensions that lead to $t$-transversals. To do so, we select certain variables (which we call core variables) that appear in the anchor and enumerate the transversals of the anchor by considering all settings (to 1/0) of the core variables depending on whether they are included in or excluded from a transversal and noting any variables which are forced as a consequence. Let $\gamma$ be one such transversal and assume that it includes $j$ variables and excludes $e$ variables in the anchor. The transversal $\gamma$ induces a restricted CNF $\mathbb{F}^{\prime}$. Specifically, $\mathbb{F}^{\prime}$ is obtained by eliminating all clauses from $\mathbb{F}$ containing any included variable of $\gamma$ and by deleting any excluded variable from each clause of $\mathbb{F}$. It is clear that $\mathbb{F}^{\prime}$ is a (tj)(t-j)-threshold monotone $3$-CNF over njen-j-e variables and its deficit is 3tn2j+e3t-n-2j+e. We also ensure that $\mathbb{F}^{\prime}\in{\Xi}_{i}$ for some $i\in\{0,1,2d,2o,3,4\}$ so we can inductively bound the number of (tj)(t-j)-transversals by $\Phi_{i}(s+k-2j,t-j)$.
We will continue with our example where $a$ always appears together with $b$. We use $\mathbb{A}$ as our anchor and $a$ and $b$ as core variables. When $a$ is set to $1$, we know $b$ must be set to 0 since if $b$ were set to $1$, $a$ would not have a critical clause. The resulting formula is of type 0 with $n^{\prime}=n-2$. Its threshold $t^{\prime}$ is at least t1t-1 and $s^{\prime}=s-1$. We now apply induction — if $s$ is even, $\Phi_{0}(s^{\prime},t^{\prime})=\Phi_{0}(s-1,t-1)=\frac{7}{18}\Phi_{0}(s,t)$, and if $s$ is odd, $\Phi_{0}(s^{\prime},t^{\prime})=\Phi_{0}(s-1,t-1)=\frac{3}{7}\Phi_{0}(s,t)$. When $a$ is set to 0, $b$ is set to either 0 or $1$. In either case, the resulting formula is of type 0. The first case results in exactly the same bounds as the case when $a$ is set to $1$. In the second case, we have to set $c$, $d$ and $e$ to $1$ so we get $n^{\prime}=n-5$, $t^{\prime}\geq t-3$ and s=s4s=s-4. We apply induction to get that the ratio of the number of transversals of the restricted formula to $\Phi_{0}(s,t)$ is at most $\frac{1}{12}$ regardless of the parity of $s$. The sum of the fractions is at most $1$, so our induction is successful.
@@
Our proof is presented in the form of a table as shown below. The columns with the label Core variables contain entries indicating whether the core variables are included in or excluded from the transversals of the anchor, $1$ signaling inclusion and 0 exclusion. The column labeled Forced variables lists variables which are forced to be included or excluded as a consequence of the setting of a subset of core variables. The columns under the label Restricted CNF parameters describe the values of the parameters and the type of the restricted CNF. Each row (except the last row) represents a transversal of the anchor. For each such row, the columns labeled Inductive bound contain the ratio of the inductive upper bounds to the claimed upper bound depending on the parity of $s$.
The last row contains the sum of the fractions. The induction step is successful if the fractions add up to a total less than or equal to $1$.
Core variables Forced variables Restricted CNF parameters Inductive bound